OpenStudy (anonymous):

How long is a string from the top of a 8ft pole to a point 5 ft from the base of the pole?

OpenStudy (amistre64):

8^2 + 5^2 = 64 + 25 = 89 sqrt(89).. simplyfy if need be

OpenStudy (anonymous):

sqrt89

OpenStudy (anonymous):

A Trig solution: \[\text{StringLength}=8\text{Csc}\left[\text{ArcTan}\left[\frac{8}{5}\right]\right]=\sqrt{89} \]